A line passing through (-5,0) and (8,y) is perpendicular to a line with a slope 13/3. Find the value of y. My head rattles when I try to work this problem! Please help. Thanks!
Answer by misscrt(37)

A line passing through (-5,0) and (8,y) is perpendicular to a line with a slope 13/3. Find the value of y. My head rattles when I try to work this problem! Please help. Thanks!
If the line has slope 13/3, then the line passing thru it that is perpendicular has a value of slope that is the negative reciprocal of 13/3, which is -3/13.
Use the formula for slope to find the value of y:
m=y2-y1/x2-x1
-3/13 =(y-0)/(8-(-5))
-3/13=y/13
Cross multiply
(-3)(13)=(13)(y)
-39=13y
Solve for y
-39/13=13y/13
-3 = y
Hope that helps!
Greetings,
misscrt